# From Movement Primitives to Distance Fields to Dynamical Systems

### v1

@@ Abstract
Developing autonomous robots capable of learning and reproducing complex motions from demonstrations remains a fundamental challenge in robotics. On the one hand, movement primitives (MPs) provide a compact and modular representation of continuous trajectories. On the other hand, autonomous systems provide control policies that are time independent. We propose in this paper a simple and flexible approach that gathers the advantages of both representations by transforming MPs into autonomous systems. The key idea is to transform the explicit representation of a trajectory as an implicit shape encoded as a distance field. This conversion from a time-dependent motion to a spatial representation enables the definition of an autonomous dynamical system with modular reactions to perturbation. Asymptotic stability guarantees are provided by using Bernstein basis functions in the MPs, representing trajectories as concatenated quadratic Bézier curves, which provide an analytical method for computing distance fields. This approach bridges conventional MPs with distance fields, ensuring smooth and precise motion encoding, while maintaining a continuous spatial representation. By simply leveraging the analytic gradients of the curve and its distance field, a stable dynamical system can be computed to reproduce the demonstrated trajectories while handling perturbations, without requiring a model of the dynamical system to be estimated. Numerical simulations and real-world robotic experiments validate our method’s ability to encode complex motion patterns while ensuring trajectory stability, together with the flexibility of designing the desired reaction to perturbations. An interactive project page demonstrating our approach is available at https://mp-df-ds.github.io/.
@@ II Representing Trajectories as Distance Fields
Given a reference trajectory $\mathbf{F}$ whose points are given by $\mathbf{f}(t):[0,T]\to\mathbb{R}^{D}$ parameterized by $t\in[0,T]$, where $T$ is the trajectory length and $D$ is the spatial dimension, the distance field $d(\mathbf{x})$ at any point $\mathbf{x}\in\mathbb{R}^{d}$ is defined as:
@@ II-A Quadratic spline as Movement Primitives
Splines are widely used as movement primitives due to their ability to provide smooth and flexible trajectory representations. We use here the term quadratic spline to define the concatenation of quadratic Bézier curves, corresponding to piecewise Bernstein polynomial functions that ensures continuity in both position and velocity. Each segment of a quadratic spline is defined over an interval $[0,1]$ and is parameterized as a weighted sum of basis functions
$\mathbf{f}_{i}(t)=(1-t)^{2}\,\mathbf{w}^{1}_{i}+2(1-t)t\,\mathbf{w}^{2}_{i}+t^{2}\,\mathbf{w}^{3}_{i},\quad t\in[0,1],$ (3)
where $\mathbf{w}^{1}_{i}$, $\mathbf{w}^{2}_{i}$, and $\mathbf{w}^{3}_{i}$ are the control points of the $i$-th segment. Quadratic splines can approximate complex trajectories by decomposing the motion into localized polynomial segments. To ensure smooth concatenation between consecutive segments, continuity constraints are imposed with
@@
where $\mathbf{f}(t)$ represents the concatenated quadratic splines, $\mathbf{T}(t)=[1,t,t^{2}]$ encodes the quadratic polynomial basis, $\mathbf{B}$ is the corresponding coefficient matrix, and $\mathbf{C}$ enforces the continuity constraints. A detailed derivation of this formulation can be found in [18].
The parameters $\mathbf{w}$ determine the shape of the quadratic spline and are estimated to best fit a given reference trajectory. Given a set of trajectory points $\{\mathbf{p}_{j}\}_{j=1}^{M}$ sampled at time instances $\{t_{j}\}_{j=1}^{M}$, we formulate the parameter estimation as an optimization problem that minimizes the discrepancy between the spline representation and the reference trajectory $\bm{f}(t)$. Specifically, we solve
$\min_{\mathbf{w}}\sum_{j=M}^{N}\left\|\mathbf{f}(t_{j})-\mathbf{p}_{j}\right\|^{2},$ (6)
@@ II-B Distance Fields of Quadratic Splines
By minimizing $c(t)=\frac{1}{2}(\mathbf{f}_{i}(t)-\mathbf{x})^{\scriptscriptstyle\top}(\mathbf{f}_{i}(t)-\mathbf{x})$ w.r.t. $t$, we get the closest point $\mathbf{f}_{i}(t)$ to $\mathbf{x}$ on the curve. By differentiating $c(t)$ and equating to zero, we get
$(\mathbf{f}_{i}(t)-\mathbf{x})^{\top}\;\mathbf{\dot{f}}_{i}(t)=0,$ (9)
where $\mathbf{\dot{f}}_{i}(t)$ denotes the derivative of the spline with respect to $t$. We can observe in the above equation that the closest point on the curve satisfies the orthogonality condition between the residual vector $\mathbf{f}_{i}(t)-\mathbf{x}$ and the tangent vector of the spline $\mathbf{\dot{f}}_{i}(t)$.
Substituting the quadratic spline formulation
$\mathbf{f}_{i}(t)$ $=(1-t)^{2}\,\mathbf{w}^{1}_{i}+2(1-t)t\,\mathbf{w}^{2}_{i}+t^{2}\,\mathbf{w}^{3}_{i}$ (10)
$\mathbf{\dot{f}}_{i}(t)$ $=-2(1-t)\,\mathbf{w}^{1}_{i}+(2-4t)\,\mathbf{w}^{2}_{i}+2t\,\mathbf{w}^{3}_{i},$
@@
$\bm{\alpha^{3}}_{i}\,t^{3}+\bm{\alpha^{2}}_{i}\,t^{2}+\bm{\alpha^{1}}_{i}\,t+\bm{\alpha^{0}}_{i}=0,$ (11)
@@
$d_{i}(\mathbf{x})=\min\big{\{}\|\mathbf{f}_{i}(t^{*})-\mathbf{x}\|:t^{*}\in[0,1]\big{\}}.$ (12)
If no roots satisfy $t^{*}\in[0,1]$, the closest point occurs at one of the segment boundaries, $t=0$ or $t=1$. The overall minimal distance to a trajectory composed of multiple segments is given by
$d(\mathbf{x})=\min\big{\{}d_{i}(\mathbf{x}):i=1,\dots,T\big{\}},$ (13)
where $d_{i}$ is the minimal distance for the $i$-th segment, and $T$ is the total number of segments.
@@ II-D From Distance Fields to Dynamical Systems
In this section, we demonstrate how the distance field representation can be utilized to design stable dynamical systems in a structured and efficient manner. Given a distance field $d(\mathbf{x})$, which represents the shortest distance from a point $\mathbf{x}$ to a reference trajectory encoded by quadratic splines, we define the dynamical system
$\dot{\mathbf{x}}=-\alpha\;\mathbf{\nabla d}(\mathbf{x})+\beta\;\dot{\mathbf{x}}_{\text{proj}},$ (15)
where $\mathbf{\nabla d}(\mathbf{x})$ is the gradient of the distance field, pointing away from the trajectory. $\mathbf{x}_{\text{proj}}$ is the closest point on the trajectory, computed using gradient projection (2). $\dot{\mathbf{x}}_{\text{proj}}$ represents the velocity along the trajectory, obtained by the analytic derivatives of the quadratic spline. $\alpha\geqslant 0$, $\beta\geqslant 0$ are gain parameters that regulate attraction to the trajectory and movement along it, respectively. The first term in the system ensures attraction toward the trajectory, while the second term facilitates smooth movement along it. To analyze the stability of this system, we introduce the following Lyapunov function
@@
$=d(\mathbf{x})\mathbf{\nabla d}(\mathbf{x})^{\top}(-\alpha\ \mathbf{\nabla d}(\mathbf{x})+\beta\ \dot{\mathbf{x}}_{\text{proj}})$
$=-\alpha\ d(\mathbf{x})\|\mathbf{\nabla d}(\mathbf{x})\|^{2}+\beta\ d(\mathbf{x})\mathbf{\nabla d}(\mathbf{x})^{\top}\dot{\mathbf{x}}_{\text{proj}}$
$=-\alpha\ d(\mathbf{x})+\beta\ d(\mathbf{x})\mathbf{\nabla d}(\mathbf{x})^{\top}\dot{\mathbf{x}}_{\text{proj}}.$
Since $d(\mathbf{x})\geq 0$, the first term in $\dot{V}(\mathbf{x})$ is always non-positive. To ensure $\dot{V}(\mathbf{x})\leq 0$, we analyze the second term $d(\mathbf{x})\mathbf{\nabla d}(\mathbf{x})^{\top}\dot{\mathbf{x}}_{\text{proj}}$ in two cases:
Case 1: Interior Point on the Trajectory. According to (9), if $\mathbf{x}_{\text{proj}}$ lies in the interior of the trajectory, the gradient $\mathbf{\nabla d}(\mathbf{x})$ is orthogonal to the velocity:
$\mathbf{\nabla d}(\mathbf{x})^{\top}\dot{\mathbf{x}}_{\text{proj}}=0.$ (18)
Case 2: Boundary Point on the Trajectory. If $\mathbf{x}_{\text{proj}}$ is at the trajectory endpoint, we have $\dot{\mathbf{x}}_{\text{proj},T}=0$. If it is at the start of the trajectory, the velocity satisfies:
$\mathbf{\nabla d}(\mathbf{x})^{\top}\dot{\mathbf{x}}_{\text{proj},1}\leq 0.$ (19)
Otherwise, the movement along the trajectory would decrease $d(\mathbf{x})$, contradicting the distance field definition. Therefore, in all cases, $\dot{V}(\mathbf{x})\leq 0$, ensuring global stability. Furthermore, $\dot{V}(\mathbf{x})=0$ only when $d(\mathbf{x})=0$, meaning the system converges asymptotically to the trajectory.
@@
where $\lambda>0$ is a parameter to balance the two terms. In essence, larger values of the distance field $d(\mathbf{x})$ correspond to regions farther from the trajectory, prompting the dynamical system to exert a stronger influence to steer the state back toward the trajectory. Conversely, smaller distance values indicate proximity to the trajectory, causing the system to reduce its corrective force and allow for smoother motion along the path.
@@ III Experiments
Figure 4: Computation time in terms of various numbers of segments and dimensions.
@@ III-A Numerical Experiments
Computation Time. We evaluate the computational cost of our approach across varying numbers of segments ($N$) and dimensions ($D$). Specifically, we measure the time required to compute distances, gradients, and time phases for 2,500 points relative to a quadratic curve in parallel using PyTorch. Results are shown in Figure 4. Notably, for typical robotic settings (e.g., $N\leq 20$, $D\leq 10$), our method is highly efficient—requiring only $\sim$2 ms on a single NVIDIA GeForce RTX 3060 laptop GPU and less than 10  ms on an AMD Ryzen 7 6800H CPU. Even with significantly larger $N$ and $D$, the method remains efficient, particularly on the GPU. Further performance gains are possible due to the method’s simplicity and ease of implementation.
@@
Figure 6: Dynamical systems using either numerical or analytical distance fields. The left plots show the produced trajectory. The right plots show the instability caused by discontinuities in the numerical gradient, learning to an unstable dynamical system.

### v2

@@ Abstract
Developing autonomous robots capable of learning and reproducing complex motions from demonstrations remains a fundamental challenge in robotics. On the one hand, movement primitives (MPs) provide a compact and modular representation of continuous trajectories. On the other hand, autonomous systems provide control policies that are time independent. We propose in this paper a simple and flexible approach that gathers the advantages of both representations by transforming MPs into autonomous systems. The key idea is to transform the explicit representation of a trajectory as an implicit shape encoded as a distance field. This conversion from a time-dependent motion to a spatial representation enables the definition of an autonomous dynamical system with modular reactions to perturbation. Asymptotic stability guarantees are provided by using Bernstein basis functions in the MPs, representing trajectories as concatenated quadratic Bézier curves, which provide an analytical method for computing distance fields. This approach bridges conventional MPs with distance fields, ensuring smooth and precise motion encoding, while maintaining a continuous spatial representation. By simply leveraging the analytic gradients of the curve and its distance field, a stable dynamical system can be computed to reproduce the demonstrated trajectories while handling perturbations, without requiring a model of the dynamical system to be estimated. Numerical simulations and real-world robotic experiments validate our method’s ability to encode complex motion patterns while ensuring trajectory stability, together with the flexibility of designing the desired reaction to perturbations.
An interactive project page demonstrating our approach is available at https://idiap.github.io/mp-df-ds/.
@@ II Representing Trajectories as Distance Fields
Given a reference trajectory $\mathbf{F}$ whose points are given by $\mathbf{f}(t):[0,T]\to\mathbb{R}^{D}$ parameterized by $t\in[0,T]$, where $T$ is the trajectory length and $D$ is the spatial dimension, the distance field $d(\mathbf{x})$ at any point $\mathbf{x}\in\mathbb{R}^{{\color[rgb]{0,0,0}D}}$ is defined as:
@@ II-A Quadratic spline as Movement Primitives
Splines are widely used as movement primitives due to their ability to provide smooth and flexible trajectory representations. We use here the term quadratic spline to define the concatenation of quadratic Bézier curves, corresponding to piecewise Bernstein polynomial functions that ensures continuity in both position and velocity. A quadratic spline is composed of $N$ segments. Each segment $\mathbf{f}_{i}$ is a quadratic Bézier curve parameterized by a local time $t_{i}\in[0,1]$:
$\mathbf{f}_{i}({\color[rgb]{0,0,0}t_{i}})=(1-{\color[rgb]{0,0,0}t_{i}})^{2},\mathbf{w}^{1}_{i}+2(1-{\color[rgb]{0,0,0}t_{i}}){\color[rgb]{0,0,0}t_{i}},\mathbf{w}^{2}_{i}+{\color[rgb]{0,0,0}t_{i}}^{2},\mathbf{w}^{3}_{i},\quad{\color[rgb]{0,0,0}t_{i}}\in[0,1],$ (3)
where $\mathbf{w}^{1}_{i}$, $\mathbf{w}^{2}_{i}$, and $\mathbf{w}^{3}_{i}$ are the control points of the $i$-th segment. The full trajectory $\mathbf{f}(t)$ is formed by mapping a global time $t\in[0,T]$ to the correct segment $i$ and its local time parameter $t_{i}$. This is achieved by assuming each of the $N$ segments has an equal duration, allowing the global time $t$ to be linearly scaled to identify the active segment and its corresponding local time in the [0,1] interval. Quadratic splines can approximate complex trajectories by decomposing the motion into localized polynomial segments. To ensure smooth concatenation between consecutive segments, continuity constraints are imposed with
@@
where $\mathbf{f}(t)$ represents the concatenated quadratic splines, $\mathbf{T}(t)=[1,t,t^{2}]$ encodes the quadratic polynomial basis, $\mathbf{B}$ is the corresponding coefficient matrix, and $\mathbf{C}$ enforces the continuity constraints. The basis function matrix $\bm{\phi}(t)$ is a piecewise function of the global time $t$ that internally handles the selection and time-scaling for the appropriate segment. A detailed derivation of this formulation can be found in [18].
The parameters $\mathbf{w}$ determine the shape of the quadratic spline and are estimated to best fit a given reference trajectory. Given a set of trajectory points $\{\mathbf{p}_{k}\}_{k=1}^{K}$ sampled at time instances $\{t_{k}\}_{k=1}^{K}$, we formulate the parameter estimation as an optimization problem that minimizes the discrepancy between the spline representation and the reference trajectory $\bm{f}(t)$. Specifically, we solve
$\min_{\mathbf{w}}\sum_{k=1}^{K}\left\|\mathbf{f}(t_{k})-\mathbf{p}_{k}\right\|^{2},$ (6)
@@ II-B Distance Fields of Quadratic Splines
By minimizing $c(t_{i})=\frac{1}{2}(\mathbf{f}_{i}(t_{i})-\mathbf{x})^{\scriptscriptstyle\top}(\mathbf{f}_{i}(t_{i})-\mathbf{x})$ w.r.t. $t_{i}$, we get the closest point $\mathbf{f}_{i}(t_{i})$ to $\mathbf{x}$ on the curve. By differentiating $c(t_{i})$ and equating to zero, we get
$(\mathbf{f}_{i}(t_{i})-\mathbf{x})^{\top}\;\mathbf{\dot{f}}_{i}(t_{i})=0,$ (9)
where $\mathbf{\dot{f}}_{i}(t_{i})$ denotes the derivative of the spline with respect to $t_{i}$. We can observe in the above equation that the closest point on the curve satisfies the orthogonality condition between the residual vector $\mathbf{f}_{i}(t_{i})-\mathbf{x}$ and the tangent vector of the spline $\mathbf{\dot{f}}_{i}(t_{i})$. Substituting the quadratic spline formulation
$\mathbf{f}_{i}(t_{i})$ $=(1-t_{i})^{2}\,\mathbf{w}^{1}_{i}+2(1-t_{i})t_{i}\,\mathbf{w}^{2}_{i}+t_{i}^{2}\,\mathbf{w}^{3}_{i}$ (10)
$\mathbf{\dot{f}}_{i}(t_{i})$ $=-2(1-t_{i})\,\mathbf{w}^{1}_{i}+(2-t_{i})\,\mathbf{w}^{2}_{i}+t_{i}\,\mathbf{w}^{3}_{i},$
@@
$\bm{\alpha^{3}}_{i}\,t_{i}^{3}+\bm{\alpha^{2}}_{i}\,t_{i}^{2}+\bm{\alpha^{1}}_{i}\,t_{i}+\bm{\alpha^{0}}_{i}=0,$ (11)
@@
$d_{i}(\mathbf{x})=\min\big{\{}\|\mathbf{f}_{i}(t_{i}^{*})-\mathbf{x}\|:t_{i}^{*}\in[0,1]\big{\}}.$ (12)
If no roots satisfy $t_{i}^{*}\in[0,1]$, the closest point occurs at one of the segment boundaries, $t_{i}=0$ or $t_{i}=1$. The overall minimal distance to a trajectory composed of multiple segments is given by
$d(\mathbf{x})=\min\big{\{}d_{i}(\mathbf{x}):i=1,\dots,{\color[rgb]{0,0,0}N}\big{\}},$ (13)
where $d_{i}$ is the minimal distance for the $i$-th segment and N is the toal number of segments.
@@ II-D From Distance Fields to Dynamical Systems
In this section, we demonstrate how the distance field representation can be utilized to design stable dynamical systems in a structured and efficient manner. Given a reference trajectory encoded by quadratic splines, we use the approach described in Section II-B to calculate the distance field $d(\mathbf{x})$, the closest point $\mathbf{x}_{\text{proj}}$ on the trajectory, and the corresponding phase variable $t_{\text{phase}}=\mathbf{f}^{-1}(\mathbf{x}_{\text{proj}})$. We define the dynamical system
$\dot{\mathbf{x}}=-\alpha\;\mathbf{\nabla d}(\mathbf{x})+\beta\;{\color[rgb]{0,0,0}\dot{\mathbf{f}}(t_{\text{phase}})},$ (15)
where $\mathbf{\nabla d}(\mathbf{x})$ is the gradient of the distance field, pointing away from the trajectory. The term $\dot{\mathbf{f}}(t_{\text{phase}})$ is the velocity vector of the reference trajectory evaluated at $t_{\text{phase}}$, obtained from the analytic derivatives of the quadratic spline. $\alpha>0$, $\beta>0$ are gain parameters that regulate attraction to the trajectory and movement along it, respectively. The first term in the system ensures attraction toward the reference trajectory (if a perturbation occurs), while the second term produces smooth movement along the reference trajectory. To analyze the stability of this system, we introduce the following Lyapunov function
@@
$=d(\mathbf{x})\mathbf{\nabla d}(\mathbf{x})^{\top}(-\alpha\ \mathbf{\nabla d}(\mathbf{x})+\beta\ {\color[rgb]{0,0,0}\dot{\mathbf{f}}(t_{\text{phase}})})$
$=-\alpha\ d(\mathbf{x})\|\mathbf{\nabla d}(\mathbf{x})\|^{2}+\beta\ d(\mathbf{x})\mathbf{\nabla d}(\mathbf{x})^{\top}{\color[rgb]{0,0,0}\dot{\mathbf{f}}(t_{\text{phase}})}$
$=-\alpha\ d(\mathbf{x})+\beta\ d(\mathbf{x})\mathbf{\nabla d}(\mathbf{x})^{\top}{\color[rgb]{0,0,0}\dot{\mathbf{f}}(t_{\text{phase}})}.$
Since $d(\mathbf{x})\geq 0$, the first term in $\dot{V}(\mathbf{x})$ is always non-positive. To ensure $\dot{V}(\mathbf{x})\leq 0$, we analyze the second term $d(\mathbf{x})\mathbf{\nabla d}(\mathbf{x})^{\top}{\color[rgb]{0,0,0}\dot{\mathbf{f}}(t_{\text{phase}})}$ in two cases:
Case 1: Interior Point ($0<t_{\text{phase}}<T$). This means the projection point $\mathbf{x}_{\text{proj}}$ lies in the interior of the trajectory, the orthogonality condition from (9) holds. This ensures that the gradient vector is perpendicular to the trajectory velocity vector, making their dot product zero:
$\mathbf{\nabla d}(\mathbf{x})^{\top}\dot{\mathbf{f}}(t_{\text{phase}})=0.$
Case 2: Boundary Points ($t_{\text{phase}}=0$ or $t_{\text{phase}}=T$).
  • At the endpoint ($t_{\text{phase}}=T$), the trajectory represented by the spline ends with zero velocity by design ($\dot{\mathbf{f}}(T)=0$), which further ensures the autonomous system halts when reaching this last point. Therefore, the dot product term $\mathbf{\nabla d}(\mathbf{x})^{\top}\dot{\mathbf{f}}(T)$ is zero.
  • At the start point ($t_{\text{phase}}=0$), we always have $\mathbf{\nabla d}(\mathbf{x})^{\top}\dot{\mathbf{f}}(0)\leq 0$. Otherwise, it would mean the trajectory initially moves into the half-space containing $\mathbf{x}$ and the start point $\mathbf{f}(0)$ is not the closest point of $\mathbf{x}$ on the entire trajectory.
Since the second term is non-positive in all cases, we confirm that $\dot{V}(\mathbf{x})\leq 0$, ensuring global stability. Furthermore, $\dot{V}(\mathbf{x})=0$ only when $d(\mathbf{x})=0$, which implies that the system asymptotically converges to the reference trajectory.
@@
where $\lambda>0$ is a parameter to balance the two terms. In essence, larger values of the distance field $d(\mathbf{x})$ correspond to regions farther from the trajectory, prompting the dynamical system to exert a stronger influence to steer the state back toward the trajectory. Conversely, smaller distance values indicate proximity to the trajectory, causing the system to reduce its corrective force and allow for smoother motion along the path. The parameters, $\alpha$ and $\beta$ in balancing attraction and progression terms can also be integrated with the phase variable $t_{\text{phase}}$ for more complex behaviors.
@@ III Experiments
Method 3 7 12 17 22
B.P. $6.18\pm 3.65$ $0.96\pm 0.60$ $0.28\pm 0.13$ $0.14\pm 0.06$ $0.10\pm 0.04$
RBF $16.51\pm 4.35$ $1.79\pm 0.61$ $0.26\pm 0.08$ $0.12\pm 0.04$ $\mathbf{0.05\pm 0.02}$
Fourier $\mathbf{5.18\pm 3.52}$ $\mathbf{0.56\pm 0.33}$ $\mathbf{0.19\pm 0.08}$ $\mathbf{0.09\pm 0.04}$ $\mathbf{0.05\pm 0.02}$
@@ III-A Numerical Experiments
Figure 4: Computation time for various numbers of segments (or demonstrations) and dimensions.
@@
Computation Time. We evaluate in Fig. 4 the computational cost of our approach across varying numbers of segments and dimensions. The computational cost of fusing $M$ demonstrations, each with $N$ segments, is equivalent to processing a single trajectory composed of $MN$ segments. We measure the time required to compute distances, gradients, and time phases for 2,500 points relative to a quadratic curve in parallel using PyTorch. Results are shown in Figure 4. Notably, for typical robotic settings (e.g., $MN\leq 100$, $D\leq 10$), our method is highly efficient—requiring less than 5ms on a single NVIDIA GeForce RTX 3060 laptop GPU and less than 10  ms on an AMD Ryzen 7 6800H CPU. Even with significantly larger $MN$ and $D$, the method remains efficient, particularly on the GPU. Further performance gains are possible due to the method’s simplicity and ease of implementation.
@@
Figure 6: Dynamical systems using either numerical or analytical distance fields. The left plots show the produced trajectory. The right plots show the instability caused by discontinuities in the numerical gradient, leading to an unstable dynamical system.
Figure 7: Qualitative comparison with baseline approaches. Black curves represent the original demonstration, while colored curves show the reproductions.
Comparison with Baselines. We compare our method against two widely-used approaches: Dynamic Movement Primitives (DMP) [2] and Stable Estimator of Dynamical Systems (SEDS) [8]. Our approach offers distinct advantages in terms of model capability and computational efficiency. Specifically, DMP and SEDS rely on a learned model of the system’s dynamics (e.g., a forcing term profile encoded with basis functions for DMP, or a Gaussian mixture model for SEDS). SEDS formulates the autonomous system as a constrained optimization problem. The constructed vector fields predominantly guide trajectories toward a single attractor, making it unsuitable for modeling and reproducing elaborate paths that would temporarily move away from the goal. It also becomes computationally challenging to apply this approach to high-dimensional spaces. In contrast, our method constructs the dynamics directly from the geometry of the reference path (without learning a model) by superposing normal attraction and tangential progression vectors. DMP is based on the advancement of a phase variable (reparameterization of time), instead of an autonomous system. It is also typically designed to learn from a single demonstration. In contrast, our approach constructs an autonomous system that naturally accommodates multiple demonstrations by fusing their respective distance fields through a simple and efficient minimum operation. Additionally, our method offers highly intuitive and modular control over the robot’s behavior. A single parameter directly balances the influence of the attraction term (how strongly we need to come back to the path if a perturbation occurs) and the progression term (how to move along the path), which can be easily balanced (see the last two plots of Fig. 7). In contrast, the parameters in DMP and SEDS are often embedded within the learned models, making them less modular.